OpenStudy (anonymous):

Solve on the interval (0,2π): 1+cosθ=1/2

OpenStudy (solomonzelman):

1+cosθ=1/2 cosθ=1/2 -1 cosθ=-1/2 

OpenStudy (solomonzelman):

cos 30 = a / 2a = 1/2 So it is -30

OpenStudy (solomonzelman):

And -30 would be the first solution, you just need to write it in your interval.

OpenStudy (anonymous):

So, the first solution would be pi/6?

OpenStudy (solomonzelman):

the first solution is -π/6

OpenStudy (solomonzelman):

as I said, because it is equal to NEGATIVE 1/2

OpenStudy (anonymous):

Ah, right, that makes sense. So would I use a reference angle to make it positive?

OpenStudy (anonymous):

Making it 11pi/6?

OpenStudy (anonymous):

Nevermind, I figured it out, thank you for your help!